OpenStudy (anonymous):

what are slop and y-intercept for the graph of -3x+=1

OpenStudy (anonymous):

rearrange it so that you get it in the form of y=mx + b so bring everything to one side

OpenStudy (anonymous):

i still dont get it. 1=3x+y ?

OpenStudy (anonymous):

okay have everything equaling y so y=-3x + 1

OpenStudy (anonymous):

wait, you changed it what is your original equation?

OpenStudy (anonymous):

-3x+y=1

OpenStudy (anonymous):

so then y=3x +1

OpenStudy (anonymous):

this equation is in the form of y=mx + b, are you familiar with that equation? Its known as the slope equation becasue m represents slope while b represents the y intercept

OpenStudy (anonymous):

*slope formula

OpenStudy (anonymous):

yes i am aware of it. so it would be y=3x+1 . therefore the y intercept would be 1 and the slope would be -3 ?

OpenStudy (anonymous):

it would be just 3 not -3 but yea

OpenStudy (anonymous):

oh ok. i get it know. solve for y then find the slope. ok thanks.

OpenStudy (anonymous):

no prob